# Logical Consistency of Large Language Models in Fact-Checking

### v1

@@ 3.2 Consistency of Complex Facts
The proof is given in the Appendix C. Since any propositional logic fact can be expressed in the DNF form, Proposition 1 encompasses any propositional logic fact. In addition, we measure the average consistency given a batch of $N$ DNF queries $Q=\left\{q_{k}=\vee_{i=1}^{n}\left(\wedge_{j=1}^{i_{m}}e^{k}_{ij}\right):1\leq k\leq N\right\}$ as $\mathtt{Consistency}(\mathtt{LLM},Q)=\frac{1}{N}\sum_{k=1}^{N}\mathds{1}_{\mathtt{LLM}(q_{k})=\vee_{i=1}^{n}\left(\wedge_{j=1}^{i_{m}}\mathtt{LLM}(e^{k}_{ij})\right)}$.
In the following, as a basic example of complex facts, we show an LLM’s consistency assessment using a conjunction with two atomic relation facts from Freebase.
@@
Consider two true facts from Freebase: $p=$ filmProduction(Universal Studios, The Family Man) and $q=$ filmProduction(Relativity Media, Valentine’s Day). By prompting Llama2222-$7$B with an LLMQuery constituting a complex fact with $\wedge$ operator, i.e., the (Fact $p$wedge Fact $q$ , Context)-pair, we obtain the response $\mathtt{LLM}(p\wedge q)=1$. Similarly, by prompting the (Fact $p$, Context)-pair, we receive the response $\mathtt{LLM}(p)=1$. We obtain the same for q, $\mathtt{LLM}(q)=1$. In this case, the LLM is consistent because $\mathtt{LLM}(p\wedge q)=\mathtt{LLM}(p)\wedge\mathtt{LLM}(q)$.
Consider the true fact $p=$ countryReverse(Brazil, Artistic gymnastics) implying that Artistic Gymnastic is played in Brazil and another true fact $q=$ sport(2012 Summer Olympics, Artistic gymnastics). Upon prompting Llama2222-$7$B with an LLMQuery with (Fact $p\wedge\text{Fact }q$ , Context)-pair, we obtain the response $\mathtt{LLM}(p\wedge q)=1$. However, upon prompting Llama2222-$7$B with the constituent simple facts individually, we receive the response $\mathtt{LLM}(p)=0$ and $\mathtt{LLM}(q)=1$. In this case, the LLM is inconsistent because $\mathtt{LLM}(p\wedge q)=1\neq 0=\mathtt{LLM}(p)\wedge\mathtt{LLM}(q)$.
@@ 3.3 Consistency for Logic Rules
An important family of true facts comes from various logical rules of inference, such as the commutative, associative, and distributive properties. One critical question is: Do LLMs respond such that these laws of inference are respected? Let $p,q,s$ be propositional logic facts. An LLM is consistent w.r.t. the commutative law if reordering logic facts results in the identical response.
@@
3.4 Constructing LLMQuery with Facts and KG Contexts
Our goal is to assess the logical consistency of LLMs on fact-checking queries from KGs, as discussed above. It is plausible that LLMs lack the necessary knowledge to answer the fact-checking query accurately, let alone being consistent, specially on recent facts never seen during pre-training/fine-tuning. As such, motivated by RAG-based framework, we intend to provide necessary knowledge or context from the KG in LLMQuery before assessing for consistency. Algorithm 1 provides a template pipeline for LLMQuery, where a key step is to find the relevant context for the target fact. Among different choices, we first discuss a graph traversal based method for context retrieval, followed by alternate methods such as vector-embedding based retrieval.
Graph Traversal. Given a target fact, we apply breadth first search (BFS) to get a relevant subgraph from the KG. For simplicity, let us consider a simple fact either in the form $r(u,v)$, or in the negation form $\neg r(u,v)$. The relevant information of an entity is generally expected to be in the local neighbourhood in the KG (Dettmers et al., 2018; Wang et al., 2020b). Hence, we conduct BFS from the subject entity $u$ up to a certain number of hops (e.g., $\lambda$ = 2 or 3) in the KG, and get a connected subgraph (line 3). To analyse the complexity of BFS, let the $\lambda$-hop neighbourhood of $u$ contain $|V_{\lambda}(u)|$ nodes and $|E_{\lambda}(u)|$ edges in the KG. The time complexity of BFS-based context building is $O(|V_{\lambda}(u)|+|E_{\lambda}(u)|)$. Thereafter, we provide the list of triplets in the subgraph as a context to LLMQuery (line 4). Our hypothesis is that the context contains necessary information for the LLM to validate the target fact: If the fact is implied by the context, the LLM is expected to return the binary response 1 (i.e., the LLM validates the fact), and 0 otherwise (i.e., the LLM invalidates it).
Optimizing Context Generation. LLMs are restricted to processing a finite length context. Besides, LLMs’ performance often deteriorates with long contexts (Liu et al., 2023a). Therefore, we further prune the context/subgraph from BFS, especially when the subject entity $u$ is connected to many other entities within a few hops. We adopt three potential approaches: (1) We prioritize triplets having semantically similar relations as the target fact’s relation $r$. We identify such semantically similar relations by utilizing the ontology and relation hierarchy, often available with KG (Chah, 2018; Chen et al., 2023). For a complex fact, we split the context size equally for each simple fact involved in it. (2) We apply vector-embedding based methods to project triplets in the subgraph to a high-dimensional embedding space  (Reimers & Gurevych, 2019) and find relevant triplets close to the embedded target fact using approximate nearest neighbor (ANN) algorithms  (Malkov & Yashunin, 2020). (3) As the last approach, we bypass applying BFS and embed the entire KG triplets and retrieve relevant triplets to the target fact. Thus, we provide the most relevant context to the LLM so that we can precisely assess its logical consistency. Furthermore, our framework for assessing consistency is modular and can benefit from future development in graph-based retrieval of related contexts (e.g., graphRAG  (Liu et al., 2024; Sanmartin, 2024)) for the LLM.
@@ 4 Improving the LLM Consistency by Supervised Fine-tuning
Figure 2: The adaptation of LLMQuery from Figure 1 for instruction prompting. While still zero-shot, it contains clear instruction (in red) to guide the LLM towards a correct response.
Supervised fine-tuning requires time and computational resources, and in some cases it is not needed at all (Radford et al., 2019; Ziegler et al., 2019). Hence, it is important to verify if our task necessitates fine-tuning. We first consider zero-shot instruction prompting – shown in Figure 2, as a computationally cheaper alternative. The intention is to instruct the LLM about when to say yes and when to say no. We instruct it to say yes when the fact triplets are inside the context, and no otherwise. The context length is limited to 1000 tokens. We empirically find that even for simple facts checking, such zero-shot instruction prompting does not improve the logical consistency, while supervised fine-tuning outperforms zero-shot instruction prompting; the consistency improves by up to 46% (see Tables 3 and 14).
The reasons for inferior performance of zero-shot instruction prompting compared to supervised fine-tuning are two-fold: (1) Nuances and intricacies of logical facts. It is well known that supervised fine-tuning of pre-trained LLMs is preferable for tasks that require precise and nuanced predictions, especially when the task involves multiple classes or fine-grained distinctions (Wei et al., 2021; Li & Liang, 2021). In such cases, providing explicit supervision through labelled data allows the LLM to adjust its behaviour based on the nuances and intricacies of the task domain and learn to make accurate predictions based on the task-specific examples (Sainz et al., 2024). For a simple fact-checking task, the intricacies arise from understanding the differences between a true fact and its corresponding false fact. (2) Limited knowledge. Fine-tuning a pre-trained LLM with task-specific labelled examples enables it to directly optimize its parameters for the target task, leading to potentially better performance compared to zero-shot instruction prompting (Ziegler et al., 2019; Xu et al., 2023), which relies on generalization from instructions given in the prompt. For instance, supervised fine-tuning has shown better performance than zero-shot prompting on tasks such as named entity recognition, relation extraction, and relation classification  (Sainz et al., 2024; Xu et al., 2023), while  Ziegler et al. (2019) showed that supervised fine-tuning with explicit task-specific instructions or preferences performs better than zero-shot prompting on text summarization task. In the case of our simple fact-checking task, without sufficient data, the LLM fails to learn significantly different representations of a false fact from a true fact.
@@ 4.2 Methodology for Supervised Fine-tuning
We consider supervised fine-tuning to align LLM responses to be logically consistent. We consider a two-steps procedure: dataset creation and supervised fine-tuning. Here, we discuss dataset creation, and refer to Appendix D for details on fine-tuning and the two-step Algorithm 2.
@@ 5 Experimental Results
  • Ablation Study. Do instruction prompting and its variants like zero-shot, few-shot, and chain-of-thought prompts improve consistency compared to fine-tuning? What is the impact of different KG retrieval methods on the consistency of the LLM? Does adding KG contexts help LLMs in improving accuracy and consistency, and how does the length of context play a role? What is the impact of hyperparameters, e.g., learning rate in fine-tuning?
@@ 5.1 Datasets and Experimental Setup
KG Datasets and Facts. We consider three KG benchmarks: Freebase (FB15151515K), NELL, and a large-scale dataset from OGB: WikiKG90909090Mv2222 (Wiki) (Hu et al., 2021b). We obtain FB15151515K and NELL from the codebase of Query2222Box (Ren et al., 2020). We create logical fact-checking datasets FreebaseLFC, NELLLFC, and WikiLFC using the FB15151515K, NELL, and Wiki KGs, respectively. Statistics of these KGs and fact-checking datasets are in Tables 4-5 in the Appendix E. In all experiments, we resort to our proposed fact-checking datasets, since – to our best knowledge – none of the earlier works focus on assessing and improving LLM’s logical consistency in fact-checking.
@@
LLMs. We evaluate the logical consistency of three open-source LLMs: the chat version of Llama2222-$7$(Touvron et al., 2023) and Llama2222-$13$(Touvron et al., 2023) with $7$B and $13$B parameters, respectively, and instruction tuned Gemma-$2$B with $2$B parameters (Team et al., 2024). We fine-tune $20$ epochs and save intermediate models. To select the best model for each dataset and fact-type, we consider the following principle: Select the optimal epoch having the maximum sum of accuracy and consistency on the evaluation set, and report results on the test set in the corresponding epoch.
@@ 5.2 Experimental results
Improved Logical Consistency via Improved Accuracy. Table 1 shows the result for simple facts $(p,\neg p)$ and facts with one logical operator. In all LLMs and FreebaseLFC dataset, fine-tuning results in a higher accuracy and consistency in most fact-types – exception is for disjunctive fact $p\vee q$ where fine-tuning demonstrates a higher accuracy yet a slight decrease in consistency (further explanation in Appendix E). We observe a similar trend in NELL and Wiki datasets – which are not included in fine-tuning – revealing the generalization effect of fine-tuning on similar facts in different datasets. Within Llama2222 base models (before fine-tuning), the $13$B model has, in general, higher accuracy and consistency on logic facts than the $7$B model – thus, more model parameters achieve improved accuracy and consistency, which is further improved via fine-tuning. Therefore, fine-tuning logic facts for increasing accuracy in one dataset increases accuracy (on average $19\%$) and logical consistency (on average $14\%$) across all datasets.
Table 2: Generalization of fine-tuning to complex facts and rules in Llama2222-$13$B. For results on Llama2222-$7$B and Gemma-$2$B, refer to Table 13 in the Appendix E
@@
Efficiency. Our empirical results demonstrate that PEFT is 3X more efficient than full fine-tuning for logical consistency. The end-to-end fact-checking time is also efficient (less than 0.63 seconds per sample) in our datasets. For details, we refer to Tables 6-7 in the Appendix E.
@@ 6 Conclusions
We propose a methodology to assess and improve the logical consistency of LLMs in fact-checking from KGs. We formalize the definition of logical consistency of LLMs on propositional logic queries comprising multiple logical operators. LLMs in general are less consistent in answering logic facts, and a fine-tuning task of recognizing the primitive logic operators such as negation, conjunction, and disjunction helps LLMs to be logically consistent – we demonstrate on average $14\%$ improvement in consistency via fine-tuning.
Our work has potential impacts on trustworthy language modelling. A logically consistent LLM is less likely to hallucinate, since its response is unlikely to change due to the logical manipulation of the prompt. Future works can be in several directions. Beyond binary LLM responses, we aim to assess consistency when LLM responses are not trivially classified into finite categories. We will extend consistency assessment to natural language queries containing logical relations. We will consider unstructured context unlike this paper, which further complicates LLM’s ability to accurately and consistently answer fact-checking queries. On the KG front, we investigate the efficacy of knowledge injection methods, or graph-to-text conversion, to input more relevant contexts in the LLMQuery. Finally, incorporating link prediction and reasoning capabilities into LLMs to conduct logic fact checking over incomplete KGs is an interesting problem.

### v2

@@ 3.2 Consistency of Complex Facts
An LLM is consistent on a DNF fact $q=\vee_{i=1}^{n}c_{i}$, where $c_{i}=\wedge_{j=1}^{i_{m}}e_{ij}$, if $\mathtt{LLM}(q)=\bigvee_{i=1}^{n}\left(\bigwedge_{j=1}^{i_{m}}\mathtt{LLM}(e_{ij})\right).$ Here, $e_{ij}$ is an atomic relation fact for any $1\leq i\leq n$ and $1\leq j\leq{i_{m}}$.
The proof is given in the Appendix C. Since any propositional logic fact can be expressed in the DNF form, Proposition 1 encompasses the consistency criteria of any propositional logic fact. In addition, we measure the average consistency given a batch of $N$ DNF queries $Q=\left\{q_{k}=\vee_{i=1}^{n}\left(\wedge_{j=1}^{i_{m}}e^{k}_{ij}\right):1\leq k\leq N\right\}$ as $\mathtt{Consistency}(\mathtt{LLM},Q)=\frac{1}{N}\sum_{k=1}^{N}\mathds{1}_{\mathtt{LLM}(q_{k})=\vee_{i=1}^{n}\left(\wedge_{j=1}^{i_{m}}\mathtt{LLM}(e^{k}_{ij})\right)}$.
The choice of DNF in Proposition 1 is not the only design choice, since the proposed method for assessing logical consistency is not reliant on any specific normal form. In fact, the definition of consistency is flexible enough to be adapted to other canonical forms such as CNF (Proposition 3). In the following, as a basic example of complex facts, we show an LLM’s consistency assessment using conjunction with two atomic relation facts from Freebase.
@@
Consider two true facts from Freebase: $p=$ $\mathtt{filmProduction}$(Universal Studios, The Family Man) and $q=$ $\mathtt{filmProduction}$(Relativity Media, Valentine’s Day). By prompting Llama2222-$7$B with an $\mathtt{LLMQuery}$ constituting a complex fact with $\wedge$ operator, i.e., the (Fact $p$wedge Fact $q$, Context)-pair, we obtain the response $\mathtt{LLM}(p\wedge q)=1$. Similarly, by prompting the (Fact $p$, Context)-pair, we receive the response $\mathtt{LLM}(p)=1$. We obtain the same for q, $\mathtt{LLM}(q)=1$. In this case, the LLM is consistent because $\mathtt{LLM}(p\wedge q)=\mathtt{LLM}(p)\wedge\mathtt{LLM}(q)$.
Consider the true fact $p=$ $\mathtt{countryReverse}$(Brazil, Artistic gymnastics) implying that Artistic Gymnastic is played in Brazil and another true fact $q=$ $\mathtt{sport}$(2012 Summer Olympics, Artistic gymnastics). Upon prompting Llama2222-$7$B with an $\mathtt{LLMQuery}$ with (Fact $p\wedge\text{Fact }q$ , Context)-pair, we obtain the response $\mathtt{LLM}(p\wedge q)=1$. However, upon prompting Llama2222-$7$B with the constituent simple facts individually, we receive the response $\mathtt{LLM}(p)=0$ and $\mathtt{LLM}(q)=1$. In this case, the LLM is inconsistent because $\mathtt{LLM}(p\wedge q)=1\neq 0=\mathtt{LLM}(p)\wedge\mathtt{LLM}(q)$.
@@ 3.3 Consistency for Logic Rules
An important family of true facts comes from various logical rules of inference, such as the commutative, associative, and distributive properties. One critical question is: Do LLMs respond such that these laws of inference are respected? Let $p,q,s$ be propositional logic facts. An LLM is consistent w.r.t. the commutative law if reordering logic facts results in an identical response.
@@
3.4 Constructing $\mathtt{LLMQuery}$ with Facts and KG Contexts
Our goal is to assess the logical consistency of LLMs on fact-checking queries from KGs, as discussed above. It is plausible that LLMs lack the necessary knowledge to answer the fact-checking query accurately, let alone being consistent, especially on recent facts never seen during pre-training/fine-tuning. As such, motivated by the RAG-based framework, we intend to provide necessary knowledge or context from the KG in $\mathtt{LLMQuery}$ before assessing for consistency. Algorithm 1 provides a template pipeline for $\mathtt{LLMQuery}$, where a key step is to find the relevant context for the target fact. Among different choices, we first discuss a graph traversal-based method for context retrieval, followed by alternate methods such as vector-embedding-based retrieval.
@@ 3.4 Constructing $\mathtt{LLMQuery}$ with Facts and KG Contexts
Graph Traversal. Given a target fact, we apply breadth-first search (BFS) to get a relevant subgraph from the KG. For simplicity, let us consider a simple fact either in the form $r(u,v)$, or in the negation form $\neg r(u,v)$. The relevant information of an entity is generally expected to be in the local neighbourhood in the KG (Dettmers et al., 2018; Wang et al., 2020b). Hence, we conduct BFS from the subject entity $u$ up to a certain number of hops (e.g., $\lambda$ = 2 or 3) in the KG, and get a connected subgraph (line 3). To analyse the complexity of BFS, let the $\lambda$-hop neighbourhood of $u$ contain $|V_{\lambda}(u)|$ nodes and $|E_{\lambda}(u)|$ edges in the KG. The time complexity of BFS-based context building is $O(|V_{\lambda}(u)|+|E_{\lambda}(u)|)$. Thereafter, we provide the list of triplets in the subgraph as a context to $\mathtt{LLMQuery}$ (line 4). Our hypothesis is that the context contains the necessary information for the LLM to validate the target fact: If the fact is implied by the context, the LLM is expected to return the binary response 1 (validating the fact), and 0 otherwise (invalidating it).
Optimizing Context Generation. LLMs are restricted to processing a finite-length context. Besides, LLMs’ performance often deteriorates with long contexts (Liu et al., 2023a). Therefore, we further prune the context/subgraph from BFS, especially when the subject entity $u$ is connected to many other entities within a few hops. We adopt three potential approaches: (1) We prioritize triplets having semantically similar relations as the target fact’s relation $r$. We identify such semantically similar relations by utilizing the ontology and relation hierarchy, often available with KG (Chah, 2018; Chen et al., 2023). For a complex fact, we split the context size equally for each simple fact involved in it. (2) We apply vector-embedding-based methods to project triplets in the subgraph to a high-dimensional embedding space (Reimers & Gurevych, 2019) and find relevant triplets close to the embedded target fact using approximate nearest neighbor (ANN) algorithms (Malkov & Yashunin, 2020). (3) As the last approach, we bypass applying BFS, embed the entire KG triplets, and retrieve relevant triplets to the target fact. Thus, we provide the most relevant context to the LLM so that we can precisely assess its logical consistency. Furthermore, our framework for assessing consistency is modular and can benefit from future development in graph-based retrieval of related contexts such as graphRAG (Liu et al., 2024; Sanmartin, 2024).
@@ 4 Improving the LLM Consistency by Supervised Fine-tuning
Figure 2: The adaptation of LLMQuery from Figure 1 for instruction prompting. The prompt contains a clear instruction (in blue) to guide the LLM towards outputting a correct response.
Supervised fine-tuning requires time and computational resources, and in some cases, it is not needed at all (Radford et al., 2019; Ziegler et al., 2019). Hence, it is important to verify if our task necessitates fine-tuning. We first consider zero-shot instruction prompting – shown in Figure 2, as a computationally cheaper alternative. Our goal is to instruct the LLM to output yes when the fact triplets are inside the context, and no otherwise. The context length is limited to 1000 tokens. We empirically find that even for simple facts checking, such zero-shot instruction prompting does not improve the logical consistency, while supervised fine-tuning outperforms zero-shot instruction prompting; the consistency improves by up to 46% (see Tables 3 and 11).
The reasons for inferior performance of zero-shot instruction prompting compared to supervised fine-tuning are two-fold: (1) Nuances and intricacies of logical facts. It is well known that supervised fine-tuning of pre-trained LLMs is preferable for tasks that require precise and nuanced predictions, especially when the task involves multiple classes or fine-grained distinctions (Wei et al., 2021; Li & Liang, 2021). In such cases, providing explicit supervision through labeled data allows the LLM to adjust its behaviour based on the nuances and intricacies of the task domain and learn to make accurate predictions based on the task-specific examples (Sainz et al., 2024). For a simple fact-checking task, the intricacies arise from understanding the differences between a true fact and its corresponding false fact. (2) Limited knowledge. Fine-tuning a pre-trained LLM with task-specific labeled examples enables it to directly optimize its parameters for the target task, leading to potentially better performance compared to zero-shot instruction prompting (Ziegler et al., 2019; Xu et al., 2023), which relies on generalization from instructions given in the prompt. For instance, supervised fine-tuning has shown better performance than zero-shot prompting on tasks such as named entity recognition, relation extraction, and relation classification  (Sainz et al., 2024; Xu et al., 2023), while  Ziegler et al. (2019) showed that supervised fine-tuning with explicit task-specific instructions or preferences performs better than zero-shot prompting on text summarization task. In the case of our simple fact-checking task, without sufficient data, the LLM fails to learn significantly different representations of a false fact from a true fact.
@@ 4.2 Methodology for Supervised Fine-tuning
We consider supervised fine-tuning to align LLM responses to be logically consistent. We consider a two-step procedure: dataset creation and supervised fine-tuning. Here, we discuss dataset creation and refer to Appendix D for details on fine-tuning and the two-step Algorithm 2.
@@ 5 Experimental Results
  • Ablation Study. Does instruction prompting improve logical consistency compared to fine-tuning across different model-sizes? What is the impact of different KG retrieval methods on the consistency of the LLM? Does adding KG contexts help LLMs in improving accuracy and consistency, and how does the length of context play a role? What is the impact of hyperparameters, e.g., learning rate in fine-tuning?
@@ 5.1 Datasets and Experimental Setup
KG Datasets and Facts. We consider three KG benchmarks: Freebase (FB15151515K), NELL, and a large-scale dataset from OGB: WikiKG90909090Mv2222 (Wiki) (Hu et al., 2021b). We obtain FB15151515K and NELL from the codebase of Query2222Box (Ren et al., 2020). We create logical fact-checking datasets FreebaseLFC, NELLLFC, and WikiLFC using the FB15151515K, NELL, and Wiki KGs, respectively. Statistics of these KGs and fact-checking datasets are in Tables 4-5 in the Appendix G. In all experiments, we resort to our proposed fact-checking datasets, since – to our best knowledge – none of the earlier works focus on assessing and improving LLM’s logical consistency in fact-checking.
@@
LLMs. We evaluate the logical consistency of three open-source LLMs: the chat version of Llama2222-$7$(Touvron et al., 2023) and Llama2222-$13$(Touvron et al., 2023) with $7$B and $13$B parameters, respectively, and instruction tuned Gemma-$2$B with $2$B parameters (Team et al., 2024). We fine-tune $20$ epochs and save the intermediate models. In order to select the best model for each dataset and fact type, we consider the following principle: we compute the sum of accuracy and consistency on both evaluation and test set (to prioritize both accuracy and consistency), find the optimal epoch having the maximum sum on the evaluation set, and report results on respective test set.
@@ 5.2 Experimental results
Improved Logical Consistency via Improved Accuracy. Table 1 shows the result for simple facts $(p,\neg p)$ and facts with one logical operator. In all LLMs and FreebaseLFC dataset, fine-tuning results in higher accuracy and consistency in most fact types – the exception is for disjunctive fact $p\vee q$ where fine-tuning demonstrates a higher accuracy yet a slight decrease in consistency (further explanation in Appendix G.3). We observe a similar trend in NELL and Wiki datasets – which are not included in fine-tuning – revealing the generalization effect of fine-tuning on similar facts in different datasets. Within Llama2222 base models (before fine-tuning), the $13$B model has, in general, higher accuracy and consistency on logic facts than the $7$B model – thus, more model parameters achieve improved accuracy and consistency, which is further improved via fine-tuning. Therefore, fine-tuning logic facts for increasing accuracy in one dataset increases accuracy (on average $19\%$) and logical consistency (on average $14\%$) across all datasets.
Table 2: Generalization of fine-tuning to complex facts and rules in Llama2222-$13$B. For results on Llama2222-$7$B and Gemma-$2$B, refer to Table 13 in the Appendix G
@@
Efficiency. Our empirical results demonstrate that PEFT is 3X more efficient than full fine-tuning for logical consistency. The end-to-end fact-checking time is also efficient (less than 0.63 seconds per sample) in our datasets. For details, we refer to Tables 6-7 in the Appendix G.
@@ 6 Conclusions
We propose a methodology to assess and improve the logical consistency of LLMs in fact-checking from KGs. We formalize the definition of logical consistency of LLMs on propositional logic queries comprising multiple logical operators. LLMs are usually less consistent in answering logic facts. A supervised fine-tuning for recognizing primitive logic operators (negation, conjunction, and disjunction) improves their logical consistency – we show a $14\%$ increase in consistency via fine-tuning.
Our work has potential impacts on trustworthy language modeling. A logically consistent LLM is less likely to hallucinate since its response is unlikely to change due to the logical manipulation of the prompt. Future works can be in several directions. Beyond binary responses, we aim to assess consistency when LLM responses are not trivially classified into finite categories. We will extend consistency assessment to natural language queries containing logical relations. We will consider unstructured context unlike this paper, which further complicates LLM’s ability to accurately and consistently answer fact-checking queries. On the KG front, we investigate the efficacy of knowledge injection methods, or graph-to-text conversion, to input more relevant contexts in the $\mathtt{LLMQuery}$. Finally, incorporating link prediction and reasoning capabilities into LLMs to conduct logic fact-checking over incomplete KGs is an interesting problem.